A total of $22,000 is to be divided between Melanie and Cameron. Melanie will receive $5,000 less then twice cameron's amount. How much will each person receive?
Answer by stanbon(75887)

A total of $22,000 is to be divided between Melanie and Cameron. Melanie will receive $5,000 less then twice cameron's amount. How much will each person receive?
------------------
Equations:
m + c = 22000
m = 2c-5000
------
Substitute for "m" and solve for "c":
2c-5000 + c = 22000
3c = $27000
c = $9000.00
---
Solve for "m"::
m = 2c-5000
m = 18000-5000
m = $13000
------------------
Cheers,
Stan H.
------------
Stan H.
----------------